# Let Occ Flow: Self-Supervised 3D Occupancy Flow Prediction

### v2

@@ Abstract
Accurate perception of the dynamic environment is a fundamental task for autonomous driving and robot systems. This paper introduces Let Occ Flow, the first self-supervised work for joint 3D occupancy and occupancy flow prediction using only camera inputs, eliminating the need for 3D annotations. Utilizing TPV for unified scene representation and deformable attention layers for feature aggregation, our approach incorporates a backward-forward temporal attention module to capture dynamic object dependencies, followed by a 3D refine module for fine-gained volumetric representation. Besides, our method extends differentiable rendering to 3D volumetric flow fields, leveraging zero-shot 2D segmentation and optical flow cues for dynamic decomposition and motion optimization. Extensive experiments on nuScenes and KITTI datasets demonstrate the competitive performance of our approach over prior state-of-the-art methods. Our project page is available at https://eliliu2233.github.io/letoccflow/.
@@ 1 Introduction
Existing occupancy flow prediciton methods [8, 9] rely heavily on detailed 3D occupancy flow annotations, which complicates their scalability to extensive training datasets. Previous research has attempted to address the dependency on 3D annotations for occupancy prediction tasks. Some strategies utilized LiDAR data to generate occupancy labels [3, 10, 11], but struggle to balance between accuracy and completeness due to LiDAR’s inherent sparsity and dynamic object movements. Recently, some methods have emerged to enable 3D occupancy predictions in a self-supervised manner by integrating the NeRF-style differentiable rendering with either weak 2D depth cues [7] or photometric supervision [5, 6]. However, existing self-supervised techniques take static world assumption as a premise and never consider the motion of foreground objects. This raises a question: Can we bridge 2D perception cues to enable self-supervised occupancy flow prediction?
To address the aforementioned issues, we introduce Let Occ Flow, a novel method for simultaneously predicting 3D occupancy and occupancy flow using only camera input, without requiring any 3D annotations for supervision, as shown in Figure 1. Our approach employs Tri-perspective View(TPV) [11] as a unified scene representation and incorporates deformable attention layers to aggregate multi-view image features. For temporal sequence inputs, we utilize a BEV-based backward-forward attention module to effectively capture long-distance dependencies of dynamic objects. Subsequently, we use a 3D refine module for spatial volume feature aggregation and transform the scene representation into fine-grained volumetric fields of SDF and flow.
Differentiable rendering techniques are then applied to jointly optimize scene geometry and object motion. Specifically, for geometry optimization, we employ a photometric consistency supervision [6]. While LiDAR data is available, we employ range rendering supervision [12, 13] for efficient geometric-aware sampling and avoiding the occlusion caused by the extrinsic parameters between LiDAR and the camera sensors. For motion optimization, we leverage semantic and optical flow cues for dynamic decomposition and supervision. These cues are convenient to obtain from pre-trained 2D models, which have shown superior zero-shot generalization capability in comparison to 3D models. Furthermore, previous studies have demonstrated that self-supervised photometric consistency [14, 15] enables the pre-trained flow estimator to effectively narrow the generalization gap across different data domains. Specifically, we employ a pre-trained 2D open-vocabulary segmentation model GroundedSAM [16] to derive a segmentation mask of movable objects. In addition, we use optical flow cues from an off-the-shelf 2D estimator [17, 18] to offer effective supervision of pixel correspondence for joint geometry and motion optimization.
@@
2. We designed a novel temporal fusion module with BEV-based backward-forward attention for efficient temporal interaction. Furthermore, we proposed a two-stage optimization strategy and dynamic disentanglement scheme to mitigate the training instability and sample imbalance problem.
@@
Figure 1: Let Occ Flow introduces a novel self-supervised training paradigm for simultaneous prediction of 3D occupancy and occupancy flow. Unlike earlier studies that relied on expensive annotations for 3D occupancy flow, our method employs readily available 2D labels to train the occupancy flow network. Additionally, it uses 2D optical flows to facilitate dynamic decomposition and optimize both scene geometry and object motion jointly.
@@ 2 Related Work
2.1 3D Occupancy Prediction
Vision-centric 3D occupancy prediction has emerged as a promising method to perceive the surrounding environment. TPVFormer [11] introduces a tri-perspective view representation from multi-camera images. SurroundOcc [3] utilizes sparse LiDAR data to generate dense occupancy ground truth as supervision. Rather than using 3D labels, RenderOcc [7] employs volume rendering to enable occupancy prediction from weak 2D semantic and depth labels. Recently, SimpleOccupancy [4], OccNeRF [5], and SelfOcc [6] propose a LiDAR-free self-supervised manner by photometric consistency. Despite the progress in static occupancy prediction, there has been comparatively less focus on occupancy flow prediction—a critical component for dynamic planning and navigation in autonomous driving.
2.2 Neural Radiance Field
Reconstructing 3D scenes from 2D images remains a significant challenge in computer vision. Among the various approaches, Neural Radiance Field (NeRF)[19] has gained increasing popularity due to its strong representational ability. To enhance training efficiency and improve rendering quality, several methods[20, 21, 22] have adopted explicit voxel-based representations. Additionally, some approaches [23, 24] introduced NeRF into surface reconstruction by developing Signed Distance Function(SDF) representation. Further advancements [25, 26, 27, 28] have been made to facilitate scene-level sparse view reconstruction by conditioning NeRF on image features, laying the groundwork for generalizable NeRF architectures.
2.3 Scene Flow
Scene flow enables autonomous systems to perceive motion in highly dynamic environments by offering dense motion estimation for each point in the scene. FlowNet3D [29] introduced deep-learning to estimate scene flow when trained solely on synthetic data. NSFP [30] optimizes a neural network to estimate the scene flow with expensive per-timestep optimization. Unlike previous methods that estimate correspondence between two input frames, some approaches proposed to combine scene flow prediction with vision-based occupancy network [8, 9], while they rely heavily on complete 3D occupancy and flow annotations. Our proposed approach addresses this gap by being the first work to jointly predict 3D occupancy and occupancy flow in a self-supervised manner, establishing a new direction for future research in this field.
@@ 3 Method
The overall framework is shown in Figure 2. In Sec.3.2, we first extract 3D volume features $V$ from multi-view temporal sequences with 2D-to-3D feature encoder. To capture the geometry and motion information, we employ an attention-based temporal fusion module and utilize a 3D refine module to ensure effective spatial-temporal interaction. Next in Sec.3.3, we construct the volumetric SDF and flow field using two MLP decoders and perform differentiable rendering to optimize the network with 2D labels in a self-supervised manner. Finally in Sec.3.4, we illustrate a flow-oriented two-stage training strategy and dynamic disentanglement scheme for joint geometry and motion optimization.
@@ 3.1 Problem Formulation
We aim to provide a unified framework for joint prediction of 3D occupancy $O_{t}$ and occupancy flow $F_{t}$ of the surrounding scenes at timestep $t$ using a temporal sequence of multi-view camera inputs. Formally, our prediction task is represented as:
@@
where $\mathcal{G}$ is a neural network, $I_{t}$ represents $N$-view images $\{I_{t}^{1},I_{t}^{2},...,I_{t}^{N}\}$ at timestamp $t$. $O_{t}\in\mathbb{R}^{H\times W\times Z}$ denotes the 3D occupancy at timestamp $t$, representing the occupied probability of each 3D volume grid with the value between 0 and 1. And $F_{t}\in\mathbb{R}^{H\times W\times Z\times 2}$ denotes the occupancy flow at timestamp $t$, representing the motion vector of each grid in the horizontal direction of 3D space.
@@
Temporal Fusion. To effectively capture temporal geometry and object motion, we design a temporal fusion module to enhance the scene representation further. Previous approaches [7, 33] utilize a residual block of convolution networks to fuse concatenated temporal volume features. However, it is hard for the convolution module to align the volume features of moving objects due to the inherent local receptive field. In contrast, the attention module has been proven to have the ability to capture long-distance dependencies [1]. Additionally, compared with voxel-based fusion, BEV-based temporal interaction is more computationally efficient without accuracy degradation [9].
@@
Specifically, our temporal fusion module integrates historical feature volume into the current frame. As depicted in Figure 3, this module comprises two main components: ego-motion alignment in voxel space and a Backward-Forward Attention Module (BFAM) in BEV space. To compensate for ego-motion, we first apply feature warping by transforming historical feature volumes to the current frame, according to the relative pose in ego coordinate:
@@ 3.3 Rendering-Based Optimization
SDF Field Regularization. Due to the explicit physical interpretation of SDF representation, regularization can be effectively applied to the volumetric SDF field to prevent over-fitting with limited supervision. Initially, an eikonal term $\mathcal{L}_{eik}=(1-\|\left\langle\nabla\mathcal{S},\bm{p}\right\rangle)\|_{2})^{2}$ is employed for each sample point along camera and LiDAR rays to promote the establishment of a valid SDF field, where $\nabla\mathcal{S}$ denotes the pre-computed gradient field of the volumetric SDF grid. Additionally, we also utilize a Hessian loss $\mathcal{L}_{H}$ on the second-order derivatives of the SDF grid, following the approach of SelfOcc [6]. Furthermore, an edge-aware smoothness loss $\mathcal{L}_{smooth}$ is employed following monodepth2 [36] to enforce neighborhood consistency of rendered depths and flows.
@@ 3.4 Flow-Oriented Optimization
Static Flow Supervision. Since optical flow cues provide the pixel correspondence between adjacent frames, they could be used to supervise the scene geometry effectively. Specifically, we utilize the dynamic mask mentioned above to decouple the static regions and then leverage the forward static flow for geometry supervision. To effectively use the abundant training data in a self-supervised manner, we sample an auxiliary previous frame based on the Gaussian distribution of the distance between the ego coordinates to offer an additional backward flow for geometric supervision.
Finally, our overall loss function is summarized as follows:
$\mathcal{L}=\mathcal{L}_{reproj}+\lambda_{flow}\mathcal{L}_{flow}+\lambda_{range}\mathcal{L}_{range}+\lambda_{eik}\mathcal{L}_{eik}+\lambda_{H}\mathcal{L}_{H}+\lambda_{smooth}\mathcal{L}_{smooth}+\lambda_{dreg}\mathcal{L}_{dreg}$ (12)
@@ 4.1 Self-supervised 3D Occupancy Prediction
4.2 Self-supervised Occupancy Flow Prediction
We conduct occupancy flow prediction on the KITTI-MOT [38] and nuScenes [42] datasets. To compare with the OccNeRF* [5] and RenderOcc* [7], we add a concatenation&conv temporal fusion module and an occupancy flow head to the original model with extra 2D optical flow labels as supervision. To compare with OccNet [8], we use the official implementation to reproduce the results with GT 3D occupancy flow labels from OpenOcc [8] as supervision.
Due to the absence of occupancy flow annotations, we report the occupancy flow evaluation metrics on the KITTI-MOT dataset [38] using LiDAR-projected depths and generated flow labels. More details are presented in the supplementary materials. On the nuScenes [42] dataset, with annotated occupancy flow ground truth, we report the RayIOU [40] and mAVE metrics as OccNet [8].
Table 2: Self-supervised occupancy flow prediction on the KITTI-MOT [38] dataset.
Method Occupancy Flow
@@ 4.2 Self-supervised Occupancy Flow Prediction
OccNeRF* [5] (w/o LiDAR) 3.547 6.986 7.976 10.445 0.204 0.272 0.306
Ours (w/o LiDAR) 3.431 3.529 6.429 7.635 0.215 0.084 0.118
OccNeRF* [5] (w/ LiDAR) 2.667 8.131 4.040 11.284 0.137 0.309 0.326
Ours (w/ LiDAR) 2.610 3.488 3.106 7.510 0.154 0.083 0.105
Table 3: Occupancy flow prediction on the nuScenes [42] dataset. For experiments using 2D labels as supervision, we use extra LiDAR ray supervision for occupancy flow prediction here.
Method Supervision Occupancy Flow
@@
OccNet [8] $\checkmark$ 29.28 39.68 50.02 39.66 1.61
OccNeRF* [5] $\checkmark$ 16.62 29.25 49.17 31.68 1.59
RenderOcc* [7] $\checkmark$ 20.27 32.68 49.92 36.67 1.63
Ours $\checkmark$ 25.49 39.66 56.30 40.48 1.45
We report the results of occupancy flow prediction in Table 2 and Table 3. Compared to other rendering-based methods [5, 7], our approach performs much better on both KITTI-MOT [38] and nuScenes [42] dataset, owing to our effective temporal fusion module and flow-oriented optimization strategy. And compared with 3D supervised OccNet [8], our approach achieves comparable performance on nuScenes [42], validating the effectiveness of our self-supervised training paradigm.
@@ 4.3 Ablation Study
BEV Pool&Forward Attn. 3.600 3.763 6.339 7.971 0.228 0.091 0.124 5.21M
@@
Temporal Fusion. We report the results of different temporal fusion settings in Table 4. Our proposed method achieves the best results. Compared to the 3D Convolution module, our method achieves better performance with lower memory cost due to our effective BEV-based backward-forward attention mechanism and the efficient temporal interaction between adjacent frames, rather than relying on global 3D convolution. Additionally, Our method effectively incorporates local and global information on the height dimension compared to the BEV-only attention module.
@@
5 Conclusion
In this study, we introduced Let Occ Flow, the first self-supervised method for joint 3D occupancy and occupancy flow prediction using only camera input. Utilizing the effective temporal fusion module and flow-oriented optimization, our approach effectively captures dynamic object dependencies, thus enhancing both scene geometry and object motion. Furthermore, our method extends the differentiable rendering to the volumetric flow field and leverages 2D flow and semantic cues for dynamic disentanglement. Consequently, our work paves the way for future research into more efficient and accurate self-supervised learning frameworks in dynamic environment perception.

### v3

@@ Abstract
Accurate perception of the dynamic environment is a fundamental task for autonomous driving and robot systems. This paper introduces Let Occ Flow, the first self-supervised work for joint 3D occupancy and occupancy flow prediction using only camera inputs, eliminating the need for 3D annotations. Utilizing TPV for unified scene representation and deformable attention layers for feature aggregation, our approach incorporates a novel attention-based temporal fusion module to capture dynamic object dependencies, followed by a 3D refine module for fine-gained volumetric representation. Besides, our method extends differentiable rendering to 3D volumetric flow fields, leveraging zero-shot 2D segmentation and optical flow cues for dynamic decomposition and motion optimization. Extensive experiments on nuScenes and KITTI datasets demonstrate the competitive performance of our approach over prior state-of-the-art methods. Our project page is available at https://eliliu2233.github.io/letoccflow/.
@@ 1 Introduction
Existing occupancy flow prediction methods [8, 9] rely heavily on detailed 3D occupancy flow annotations, which complicates their scalability to extensive training datasets. Previous research has attempted to address the dependency on 3D annotations for occupancy prediction tasks. Some strategies utilized LiDAR data to generate occupancy labels [3, 10, 11], but struggle to balance between accuracy and completeness due to LiDAR’s inherent sparsity and dynamic object movements. Recently, some methods have emerged to enable 3D occupancy predictions in a self-supervised manner by integrating the NeRF-style differentiable rendering with either weak 2D depth cues [7] or photometric supervision [5, 6]. However, existing self-supervised techniques take static world assumption as a premise and never consider the motion of foreground objects. This raises a question: Can we bridge 2D perception cues to enable self-supervised occupancy flow prediction?
To address the aforementioned issues, we introduce Let Occ Flow, a novel method for simultaneously predicting 3D occupancy and occupancy flow using only camera input, without requiring any 3D annotations for supervision, as shown in Figure 1. Our approach employs Tri-perspective View(TPV) [11] as a unified scene representation and incorporates deformable attention layers to aggregate multi-view image features. For temporal sequence inputs, we utilize an attention-based temporal fusion module to effectively capture long-distance dependencies of dynamic objects. Subsequently, we use a 3D refine module for spatial volume feature aggregation and transform the scene representation into fine-grained volumetric fields of SDF and flow.
Differentiable rendering techniques are then applied to jointly optimize scene geometry and object motion. Specifically, for geometry optimization, we employ a photometric consistency supervision [6] and optional range rendering supervision [12, 13]. For motion optimization, we leverage semantic and optical flow cues for dynamic decomposition and supervision. These cues are convenient to obtain from pre-trained 2D models, which have shown superior zero-shot generalization capability in comparison to 3D models. Furthermore, previous studies have demonstrated that self-supervised photometric consistency [14, 15] enables the pre-trained flow estimator to effectively narrow the generalization gap across different data domains. Specifically, we employ a pre-trained 2D open-vocabulary segmentation model GroundedSAM [16] to derive a segmentation mask of movable objects. In addition, we use optical flow cues from an off-the-shelf 2D estimator [17, 18] to offer effective supervision of pixel correspondence for joint geometry and motion optimization.
@@
2. We designed a novel attention-based temporal fusion module for efficient temporal interaction. Furthermore, we proposed a flow-oriented optimization strategy to mitigate the training instability and sample imbalance problem.
@@
Figure 1: Let Occ Flow introduces a novel self-supervised training paradigm for 3D occupancy and occupancy flow prediction. Unlike earlier studies that relied on expensive annotations for 3D occupancy flow, our method employs readily available 2D labels to train the occupancy flow network.
@@ 2 Related Work
3D Occupancy Prediction. Vision-based 3D occupancy prediction has emerged as a promising method to perceive the surrounding environment. TPVFormer [11] introduces a tri-perspective view representation from multi-camera images. SurroundOcc [3] utilizes sparse LiDAR data to generate dense occupancy ground truth as supervision. Rather than using 3D labels, RenderOcc [7] employs volume rendering to enable occupancy prediction from weak depth labels. Recent works [4, 5, 6] propose a LiDAR-free self-supervised manner by photometric consistency. Despite the progress in static occupancy prediction, there has been comparatively less focus on occupancy flow prediction—a critical component for dynamic planning and navigation in autonomous driving.
Neural Radiance Field. Reconstructing 3D scenes from 2D images remains a significant challenge in computer vision. Among the various approaches, Neural Radiance Field (NeRF)[19] has gained increasing popularity due to its strong representational ability. To enhance training efficiency and improve rendering quality, several methods[20, 21, 22] have adopted explicit voxel-based representations. Additionally, some approaches [23, 24] introduced NeRF into surface reconstruction by developing Signed Distance Function(SDF) representation. Further advancements [25, 26, 27, 28] have been made to facilitate scene-level sparse view reconstruction by conditioning NeRF on image features, laying the groundwork for generalizable NeRF architectures.
Dynamic Driving Scene Reconstruction. Recent approaches proposed to combine flow prediction with vision-based occupancy network [8, 9] for dynamic scene reconstruction, while they rely heavily on complete 3D occupancy and flow annotations. Previous approach [13] has successfully applied NeRF into self-supervised learning of dynamic driving scene reconstruction, while it typically requires expensive per-scene optimization. To the best of our knowledge, we are the very first work to achieve generalizable dynamic driving scene reconstruction in a self-supervised manner.
@@ 3 Method
The overall framework is shown in Figure 2. In Sec.3.2, we first extract 3D volume features $V$ from multi-view temporal sequences with 2D-to-3D feature encoder. To capture the geometry and motion information, we employ an attention-based temporal fusion module and a 3D refine module to ensure effective spatial-temporal interaction. In Sec.3.3, we construct the volumetric SDF and flow field using two MLP decoders and perform differentiable rendering to optimize the network in a self-supervised manner. Finally in Sec.3.4, we illustrate a flow-oriented two-stage training strategy and dynamic disentanglement scheme for joint geometry and motion optimization.
@@ 3.1 Problem Formulation
We aim to predict 3D occupancy $O_{t}$ and occupancy flow $F_{t}$ of the surrounding scenes at timestep $t$ using a temporal sequence of multi-view camera inputs. Formally, the task is represented as:
@@
where $\mathcal{G}$ is a neural network, $I_{t}$ represents $N$-view images $\{I_{t}^{1},I_{t}^{2},...,I_{t}^{N}\}$ at timestamp $t$. $O_{t}\in\mathbb{R}^{H\times W\times Z}$ denotes the 3D occupancy at timestamp $t$, representing the occupied probability of each 3D volume grid with the value between 0 and 1. And $F_{t}\in\mathbb{R}^{H\times W\times Z\times 2}$ denotes the occupancy flow at timestamp $t$, representing the object motion vector excluding ego-motion of each grid in the horizontal direction of 3D space, as described in OccNet [8].
@@
Temporal Fusion. To effectively capture temporal geometry and object motion, we design a temporal fusion module to enhance the scene representation further. Specifically, our temporal fusion module integrates historical feature volume into the current frame. As depicted in Figure 3, this module comprises two main components: ego-motion alignment in voxel space and a Backward-Forward Attention Module (BFAM) in BEV space. To compensate for ego-motion, we first apply feature warping by transforming historical feature volumes to the current frame, according to the relative pose in ego coordinate:
@@ 3.4 Flow-Oriented Optimization
Static Flow Supervision. Since optical flow cues provide the pixel correspondence between adjacent frames, they could be used to supervise the scene geometry effectively. Specifically, we utilize the dynamic mask mentioned above to decouple the static regions and then leverage the forward static flow for geometry supervision. To effectively use the abundant training data in self-supervised training, we sample an auxiliary previous frame based on the Gaussian distribution of the distance between the ego coordinates to offer an additional backward flow for geometric supervision.
Finally, with the SDF regularization loss $L_{eik},L_{H},L_{smooth}$, our overall loss function is as follows:
$\mathcal{L}=\mathcal{L}_{reproj}+\lambda_{flow}\mathcal{L}_{flow}+\lambda_{range}\mathcal{L}_{range}+\lambda_{dreg}\mathcal{L}_{dreg}+\lambda_{eik}\mathcal{L}_{eik}+\lambda_{H}\mathcal{L}_{H}+\lambda_{smooth}\mathcal{L}_{smooth}$ (12)
@@ 4.1 Self-supervised 3D Occupancy Prediction
4.2 Self-supervised Occupancy and Occupancy Flow Prediction
We conduct experiments on 3D occupancy and occupancy flow prediction on the KITTI-MOT [37] and nuScenes [39] datasets. We use OccNeRF [5] and RenderOcc [7] as rendering-based baselines. To compare with OccNeRF, we employ a concatenation&conv temporal fusion module as in RenderOcc. We also add a flow head to these two methods with extra 2D optical flow labels as supervision for fair comparison. To compare with OccNet [8], we use their official implementation to reproduce the results with GT 3D occupancy flow labels from OpenOcc [8] as supervision.
Due to the absence of occupancy flow annotations, we report the occupancy flow evaluation metrics on the KITTI-MOT dataset [37] using LiDAR-projected depths and generated flow labels. More details are presented in the supplementary materials. On the nuScenes [39] dataset, with annotated occupancy flow ground truth, we report the RayIoU [36] and mAVE metrics as in OccNet [8].
Table 2: Self-supervised occupancy and occupancy flow prediction on KITTI-MOT [37].
Method Supervision 3D Occupancy & Occupancy Flow
@@ 4.2 Self-supervised Occupancy and Occupancy Flow Prediction
OccNeRF-C* [5] C 3.547 6.986 7.976 10.445 0.204 0.272 0.306
Ours-C C 3.431 3.529 6.429 7.635 0.215 0.084 0.118
OccNeRF-L* [5] C+L 2.667 8.131 4.040 11.284 0.137 0.309 0.326
Ours-L C+L 2.610 3.488 3.106 7.510 0.154 0.083 0.105
Table 3: 3D occupancy and occupancy flow prediction on nuScenes [39].
Method Supervision 3D Occupancy & Occupancy Flow
@@
OccNet [8] 3D 29.28 39.68 50.02 39.66 1.61
OccNeRF-C* [5] C 9.93 19.06 35.84 21.61 1.53
Ours-C C 17.49 28.52 44.33 30.12 1.42
RenderOcc* [7] L 20.27 32.68 49.92 36.67 1.63
OccNeRF-L* [5] C+L 16.62 29.25 49.17 31.68 1.59
Ours-L C+L 25.49 39.66 56.30 40.48 1.45
We report the results in Table 2 and Table 3, C and L denote Camera and Lidar supervision. Compared to rendering-based methods [5, 7], our approach performs much better on both KITTI-MOT [37] and nuScenes [39], owing to our effective temporal fusion module and flow-oriented optimization strategy. Compared to 3D supervised OccNet [8], our approach achieves comparable performance on nuScenes [39], validating the effectiveness of our self-supervised training paradigm.
@@ 4.3 Ablation Study
BEV Pooling & Forward Attn. 3.600 3.763 6.339 7.971 0.228 0.091 0.124 5.21M
BEV Pooling & BFAM 3.462 3.583 6.409 7.673 0.211 0.089 0.122 6.62M
@@
Temporal Fusion. We report the results of different temporal fusion settings in Table 4. Compared to the 3D Convolution module in [7], our method achieves better performance with lower memory cost due to our efficient BEV-based attention mechanism and the effective long-distance temporal interaction. Moreover, our BFAM module outperforms the forward-only attention method introduced in [8] due to its enhanced temporal interaction capabilities. Additionally, Our method effectively incorporates local and global information in the height dimension compared to BEV-only attention.
@@
5 Discussion
Limitation and Future work. Although we use temporal sequence input to better exploit the historical information, our model cannot completely handle the occlusion problem due to the inherent rendering-based limitation. Subsequent research could investigate long-term occupancy flow modeling and solutions to leverage the temporal sequence supervision to scale up the visible range of perspective. In addition, the accuracy of occupancy flow prediction relies on the quality of optical flow cues. Future work can pay more attention on the improvement of the flow supervision quality. Finally, our occupancy flow prediction does not explicitly enforce consistency within instances, and future work may explore to integrate instance perception into occupancy flow prediction.
Conclusion. We introduced Let Occ Flow, the first self-supervised method for vision-based 3D occupancy and occupancy flow prediction. Utilizing the effective temporal fusion module and flow-oriented optimization, our approach effectively captures dynamic object dependencies, thus enhancing both scene geometry and object motion. Furthermore, our method extends the differentiable rendering to the volumetric flow field for self-supervised learning. Consequently, our work paves the way for future research into more efficient and accurate self-supervised learning frameworks in dynamic environment perception.
Acknowledgments
We sincerely thank the anonymous reviewers for their helpful comments in revising the paper. This work was supported in part by the National Nature Science Foundation of China under Grant 62373322, and in part by Zhejiang Provincial Natural Science Foundation of China under Grant No. LD24F030001.